# Live Fitness Coaching as a Testbed for Situated Interaction

### v1

@@ D.3 Long-range Videos
LoRA [25] is used for fine-tuning our model in this stage on QEVD-FIT-COACH with a learning rate of $1\times 10^{-6}$ and LoRA dimension as $64$. Other training details remain the same as the previous stage.
Data preparation. Due to memory constraints, we train on individual exercise segments of length $128$ video features which corresponds to roughly $30$ seconds. We leave it to future work to train across timed exercise transitions. Sequences are prepared as an interleaved sequence of <next> tokens and feedback response tokens. Feedback response tokens are prefixed with the special <feedback> action token and added in the sequence according to its ground truth timestamp. Finally, the sequence is prefixed with the following system prompt:

### v2

@@ D.3 Long-range Videos
LoRA [27] is used for fine-tuning our model in this stage on Qevd-Fit-Coach with a learning rate of $1\times 10^{-6}$ and LoRA dimension as $32$. Other training details remain the same as the previous stage.
Data preparation. Due to memory constraints, we train on individual exercise segments which corresponds to a length of roughly $30$ seconds. We leave it to future work to train across timed exercise transitions. Sequences are prepared as an interleaved sequence of <next> tokens and feedback response tokens. Feedback response tokens are prefixed with the special <feedback> action token and added in the sequence according to its ground truth timestamp. Finally, the sequence is prefixed with the following system prompt:
@@ E.3 Vison-language Baselines
E.4 Alternative Evaluations
To ensure the accuracy of the LLM-based auto-evaluation method described in Sec. 5.1, we perform the evaluation with alternative LLMs and compare them to human evaluations. Human evaluations are done on a random subset of $200$ feedbacks. As shown in Tab. 6, the overall ranking order is preserved across all evaluation methods.
Table 6: Evaluation of models fine-tuned with the Fit-Coach dataset on the Fit-Coach benchmark. (indicates results of non-interactive models evaluated at regular intervals, indicates human evaluation is conducted on a smaller set of $200$ feedbacks.)
Evaluation Method
Socratic-LLaMA-2-7B
Video-ChatGPT [42] (fine-tuned)
Stream-VLM
Stream-VLM (w/o 3D CNN)
Stream-VLM (w/o Action-Tokens)
Human 2.63 2.59 2.80 2.51 2.71
Mixtral-Instruct-0.1 [28] 2.39 2.42 2.56 2.17 2.56
LLaMA-3-8B-Instruct [3] 1.74 1.82 1.90 1.62 1.89
LLaMA-3-70B-Instruct [3] 2.17 2.33 2.45 2.11 2.41